# PRIS: Practical robust invertible network for image steganography

### v1

@@ 1 Introduction
Steganography is the art of concealing information. The goal of steganography is to hide a secret message within a host medium [28, 25, 4]. The host medium containing the hidden secret message is called the container [32]. The container is typically publicly visible, but the difference between the host and container should be invisible to third parties. There are some good introductions to steganography in [28, 27, 17]. Image steganography aims to hide message such as image, audio, and text within a host image in an undetectable way [44], meaning the host medium in image steganography is an image. Nowadays, image steganography is widely used in fields such as copyright protection, digital communication, information certification, and more [10].
@@ 2.2 Deep learning-based image steganography
Recently, many researchers have applied deep learning methods to image steganography [2, 19, 34, 45, 47, 49, 50, 52, 48, 4, 39, 46] and achieved better performance than traditional methods. HiDDeN [52] and SteganoGAN [49] utilize the encoder-decoder architecture to realize the embedding and extraction of secret messages and employ a third network to resist steganalysis adversarially. Shi [39] proposes Ssgan for image steganography, which is based on generative adversarial networks. Yang [46] take an U-Net structure as the generator to embed secret messages. Baluja [5, 4] and Zhang [48] hide a secret image of the same size as the host image with deep learning method, achieve a much higher payload capacity than traditional methods.
@@ 2.3 Invertible neural network
Dinh [12] proposed Invertible neural network (INN) in 2014. It learns a stable bijective mapping between a data distribution $p_{X}$ and a latent distribution $p_{Z}$. Unlike CycleGAN [53], which uses two generators and a cycle loss to achieve bidirectional mapping, INN performs both forward and backward propagation within the same network, acting as both feature encoder and image generator. INNs are also useful for estimating the posterior of an inverse problem [3]. More flexible INNs are built with masked convolutions under some composition rules in [40]. An unbiased flow-based generative model is proposed in [11]. Other works improve the coupling layer for density estimation, such as Glow [30], i-ResNet [7] and i-RevNet [23], resulting in better generation quality.
Lu [32] and Jing [24] introduced INN into image steganography. The strict invertibility of INN just meets the requirement that embedding and extraction are mutually inverse processes. Therefore they gain state-of-the-art performance in image steganography. However, the strict invertibility also means that when the container image is attacked, the noise is also transmitted to the extracted image, which make it vulnerable to attack. Xu [44] proposed robust invertible image steganography (RIIS) and improve the robustness of the aforementioned invertible neural networks.
@@ 4 Experiment
We train our model on the DIV2K training dataset [1] and evaluate it on the DIV2K testing dataset on RTX3090. The input images are cropped to a resolution of 224 $\times$ 224 pixels. For the training dataset, we apply random cropping for better generalization. For the testing dataset, we use center cropping to avoid randomness in evaluation. Each stage consists of 1600 epochs and we use the Adam [29] optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. The initial learning rates are $10^{-4.5}$ and $10^{-5.5}$ for steps 1, 2 and step 3 respectively, reducing them by half every 200 epochs for all steps. We set $\lambda_{c}$ and $\lambda_{s}$ to 1 unless otherwise specified.
@@ 4.1 Ablation experiments
4.2 Comparison with SOTA
@@
We evaluate our PRIS method against the state-of-the-art methods RIIS[44] and HiNet[24] on different levels. RIIS can handle level 1, 2 and 4 tasks, HiNet can only handle level 4 task since it does not consider attacks, and our PRIS can handle level 1, 3 and 4 tasks. To handle level 4 task, PRIS is trained for all attack methods with a 3-step strategy. To handle level 3 task, we take the trained PRIS model from level 4 task, and fine-tune different enhancement modules for different attack methods, i.e., the second step in the 3-step strategy. This means that in practice, we can select different enhancement modules for different attacks during the extraction process.
Table 4 presents the results of RIIS and PRIS on level 1 task. Our PRIS achieves an average improvement of 4.16 and 2.22 dB on PSNR-C and PSNR-S compared with RIIS respectively, with the maximum improvement being 8.32 and 6.58 dB on PSNR-C and PSNR-S under Gaussian 1 distortion. Fig. 7 is the visual results of PRIS in different type of attacks on level 1, the cover-pair and secret-pair are undistinguished from human eyes under different distortions.
Table 5 compares the results of RIIS* (RIIS on level 2) and PRIS (level 4). Although level 3 requires less information than level 2, which means that level 3 is more challenging than level 2, PRIS still achieved an average improvement of 1.34 dB on PSNR-S.
Table 6 shows the results of HiNet, RIIS-Blind (RIIS on level 4) and PRIS on level 4 task. PRIS outperformed RIIS-Blind by 0.74 dB on PSNR-S in average. HiNet has the highest PSNR-C because it did not consider attacks during training, and the attacks occurred after the embedding process, so they did not affect PSNR-C. However, the PSNR-S of HiNet was unacceptable, especially for Gaussian 10 and JPEG distortions.
@@
Table 5: Comparison (PSNR-C / PSNR-S) between different methods, where RIIS* and PRIS are trained and evaluated on level 2 and 3 respectively.
@@
Figure 7: Visual results of PRIS-A in different type of attacks.
@@ 4.3 Experiments in Practice
Rounding is an inevitable distortion in practice, and it is the first attack will occur once the embedding is done. Thus, we applied rounding error prior to Gaussian or JPEG distortion respectively, which is more realistic, and labelled these two errors RGaussian and RJPEG. Furthermore, since the type of distortion that occurs when embedding the secret image is unpredictable, we trained our PRIS on level 3 and 4 with these realistic distortion. Table 7 shows that the PNSR-C is 30.3 dB and the lowest PNSR-S on level 3 and 4 are 27.00 and 26.79 dB respectively, which are satisfactory. We also deployed PRIS in http://yanghang.site/hide/, where users can embed a secret image into a host image and retrieve it by clicking the mouse.
@@ 5 Conclusion
This paper presents a practical robust invertible network for image steganography, named PRIS. It incorporates two enhancement modules and a 3-step training strategy into invertible neural networks. Moreover, previous studies neglect the rounding distortion; however, it is unavoidable in practice, and its non-differentiability poses a challenge for training. Therefore, this paper proposes GAF to address this problem and applies rounding distortion before other distortions, which is more realistic. In addition, despite being trained on a resolution of 224 × 224 pixels, PRIS can be adapted to any resolution in practice. These features endow PRIS with the ability to outperform the existing SOTA method RIIS in terms of robustness and practicability.

### v2

@@ 1 Introduction
Steganography is the art of concealing information. The goal of steganography is to hide a secret message within a host medium [1, 2]. The host medium containing the hidden secret message is called the container [3]. The container is typically publicly visible, but the difference between the host and container should be invisible to third parties. There are some good introductions to steganography in [1, 4, 5]. Image steganography aims to hide message such as image, audio, and text within a host image in an undetectable way [6], meaning the host medium in image steganography is an image. Nowadays, image steganography is widely used in fields such as copyright protection, digital communication, information certification, and more [7].
@@ 2.2 Deep learning-based image steganography
Recently, many researchers have applied deep learning methods to image steganography and achieved better performance than traditional methods. HiDDeN [32] and SteganoGAN [33] utilize the encoder-decoder architecture to realize the embedding and extraction of secret messages and employ a third network to resist steganalysis adversarially. Shi [34] proposes Ssgan for image steganography, which is based on generative adversarial networks. Baluja [17, 15] and Zhang [16] hide a secret image of the same size as the host image with deep learning method, achieve a much higher payload capacity than traditional methods.
@@ 2.3 Invertible neural network
Dinh [35] proposed Invertible neural network (INN) in 2014. It learns a stable bijective mapping between a data distribution $p_{X}$ and a latent distribution $p_{Z}$. Unlike CycleGAN [36], which uses two generators and a cycle loss to achieve bidirectional mapping, INN performs both forward and backward propagation within the same network, acting as both feature encoder and image generator. INNs are also useful for estimating the posterior of an inverse problem [37]. More flexible INNs are built with masked convolutions under some composition rules in [38]. An unbiased flow-based generative model is proposed in [39]. Other works improve the coupling layer for density estimation, such as Glow [40] and i-ResNet [41], resulting in better generation quality.
Lu [3], Jing [21] and Jia [42] introduced INN into image steganography. The strict invertibility of INN just meets the requirement that embedding and extraction are mutually inverse processes. Therefore they gain state-of-the-art performance in image steganography. However, the strict invertibility also means that when the container image is attacked, the noise is also transmitted to the extracted image, which make it vulnerable to attack. Xu [6] proposed robust invertible image steganography (RIIS) and improve the robustness of the aforementioned invertible neural networks.
@@ 4 Experiment
4.1 Implementation details
We train our model on the DIV2K training dataset [45] and evaluate it on the DIV2K testing dataset on RTX3090. The input images are cropped to a resolution of 224 $\times$ 224 pixels. For the training dataset, we apply random cropping for better generalization. For the testing dataset, we use center cropping to avoid randomness in evaluation. Each stage consists of 1600 epochs and we use the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. The initial learning rates are $10^{-4.5}$ and $10^{-5.5}$ for steps 1, 2 and step 3 respectively, reducing them by half every 200 epochs for all steps. We set $\lambda_{c}$ and $\lambda_{s}$ to 1 unless otherwise specified.
@@ 4.2 Ablation experiments
4.3 Comparison with other methods
@@
We evaluate our PRIS method against several methods on different levels. To handle level 4 task, PRIS is trained for all attack methods with a 3-step strategy. To handle level 3 task, we take the trained PRIS model from level 4 task, and finetune different enhancement modules for different attack methods, i.e., the second step in the 3-step strategy. This means that in practice, we can select different enhancement modules for different attacks during the extraction process.
Table 4 displays the performance results of PRIS and several other methods on the level 1 task. Our PRIS exhibits superior performance when compared to the latest methods, establishing it as the most robust model. Fig. 7 is the visual results of PRIS in different type of attacks on level 1, the cover-pair and secret-pair are undistinguished from human eyes under different distortions.
@@
Table 5 compares the results of RIIS (RIIS on level 2) and PRIS (level 4). Although level 3 requires less information than level 2, which means that level 3 is more challenging than level 2, PRIS still performance better than RIIS.
Table 5: Comparison (PSNR-C / PSNR-S) between different methods, where RIIS and PRIS are trained and evaluated on level 2 and 3 respectively.
@@
Table 6 shows the results of HiNet, RIIS and PRIS on level 4 task. PRIS outperformed RIIS by 0.74 dB on PSNR-S in average. HiNet has the highest PSNR-C because it did not consider attacks during training, and the attacks occurred after the embedding process, so they did not affect PSNR-C. However, the PSNR-S of HiNet was unacceptable, especially for Gaussian $\sigma=10$ and JPEG distortions. Which indicates HiNet is fragile to attacks.
@@
Figure 7: Visual results of PRIS in different type of attacks.
@@ 4.4 Experiments in Practice
Rounding is an inevitable distortion in practice, and it is the first attack will occur once the embedding is done. Thus, we applied rounding error prior to Gaussian or JPEG distortion respectively, which is more realistic, and labelled these two errors RGaussian and RJPEG. Furthermore, since the type of distortion that occurs when embedding the secret image is unpredictable, we trained our PRIS on level 3 and 4 with these realistic distortion. Table 7 shows that the PNSR-C is 30.3 dB and the lowest PNSR-S on level 3 and 4 are 27.00 and 26.79 dB respectively, which are satisfactory. We also deployed PRIS in http://yanghang.site/hide/, where users can embed a secret image into a host image and retrieve it by simply clicking the mouse.
@@ 5 Conclusion
This paper presents a practical robust invertible network for image steganography, named PRIS. It incorporates two enhancement modules and a 3-step training strategy into invertible neural networks. Moreover, previous studies neglect the rounding distortion; however, it is unavoidable in practice, and its non-differentiability poses a challenge for training. Therefore, this paper proposes GAF to address this problem and applies rounding distortion before other distortions, which is more realistic. In addition, despite being trained on a resolution of 224 $\times$ 224 pixels, PRIS can be adapted to any resolution in practice. These features endow PRIS with the ability to outperform the existing SOTA method RIIS in terms of robustness and practicability.